OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

to get x intercept put y=0 (since x intercept is on x axis) solve for x .do the same for y

OpenStudy (anonymous):

x-intercept: (2, 0), y-intercept: (0, −3) (my anwser)

OpenStudy (radar):

Arrange equation to the "slope intercept" form. 2y = 3x - 6 Solve for y with x=0 2y=-6 y=-3 (0,-3) Now solve for x with y=0 0=3x-6 3x=6 x=2 (2,0)

OpenStudy (radar):

A very likeable answer.